### v1

@@ 1 Introduction
Link prediction in bipartite networks, or bipartite link prediction, is an important task for finding potential relations between two groups of entities. A bipartite network comprises two disjoint sets of entities and a set of edges, each connecting only two nodes from different sets. Two entities connected with an edge are considered related [1, 2]. Real-world bipartite networks often have missing or unobserved links, giving rise to the link prediction task for predicting the absence or presence of these links [3, 4, 5]. Bipartite link prediction task has practical applications such as product recommendation [6, 7], chemical-disease interactions prediction [8, 9, 10] and friend suggestion in social networking services (SNS) [2, 1, 5].
@@
Most bipartite link prediction methods rely on structural features of the networks [11]. Among all structural features, recent research highlights the importance of bi-cliques on bipartite link prediction [11, 12, 13, 14]. A bi-clique is a complete sub-graph of a bipartite network where each entity in one set is connected to each entity in the other set. Bi-cliques can be regarded as clusters of strongly interconnected entities, representing cohesive groups within networks. For example, in a chemical-disease network, a bi-clique represents a group of similar chemicals and their affected diseases. Previous research has proposed the structural hole theory [15], suggesting that two strongly connected clusters sharing numerous nodes likely belong to a larger cluster, implying that current unlinked pairs within these clusters are likely to be connected. Thus, link prediction can be performed by extracting bi-cliques and analyzing their relationships. Various methods utilizing bi-cliques have been proposed and shown good performance in small datasets.
@@
However, all them face a severe problem: they are inapplicable to large-scales datasets, as they need to extract all bi-cliques from the network, which requires significant computational resources. Such an exhaustive enumeration is usually conducted by converting the network into an equivalent formal context and performing a full process of formal concept analysis (FCA) to extract all formal concepts, each corresponding to a maximal bi-clique in the original bipartite network, from the converted formal context. The time complexity of the full process of FCA is $O(C)$ where $C$ is the number of formal concepts, i.e., which is exponential to the number of entities and links in the network. Therefore, this bi-cliques extraction step may take a long time. Furthermore, after extraction, some methods [16, 17] require training complex neural-network models from all extracted bi-cliques, consuming extensive RAM and GPU memory. Consequently, these methods struggle with large and dense bipartite networks due to their high computational cost.
@@
To overcome this problem, we propose a novel and efficient method called BicliqueEncoder, which can be applied to large datasets with manageable and adjustable memory usage and execution time. The core feature of this method is that it does not extract all bi-cliques, but only extracts those that are considered “significant” in terms of their size. As introduced above, the process of extracting bi-cliques is equivalent to that of extracting formal concepts from a formal context converted from the original bipartite network. In the field of FCA, researchers have proposed the idea of iceberg concept lattice which figures out the possibility of only extracting the formal concepts of large sizes [18] without needing to exhaustively enumerate all formal concepts. Since large formal concepts correspond to large bi-cliques which contain more information on the relations between entities, we consider they might be more significant and contribute more to link prediction than smaller bi-cliques. Hence, if we only extract and process the bi-cliques corresponding to the formal concepts in the iceberg concept lattice, we may be able to capture the key information provided by the bi-cliques with a lot fewer computational resources.
@@
We adopt the Transformer encoder [19] to learn useful information and make link predictions because we believe that it can counterbalance the information loss caused by the extracted bi-cliques. Certainly, by applying the strategy that only extracts “significant” bi-cliques, we will leave out the information of all other “insignificant” bi-cliques in the bipartite network, which may more or less cause a negative influence on the final prediction performance. To counterbalance such negative influence, we use the Transformer encoder to effectively process and capture the information from the original bipartite network as well as the extracted significant bi-cliques, and use the captured information to make link predictions. Transformer is a sequence-to-sequence deep learning framework adept at learning the dependencies within sequential data. Since bi-cliques are equivalent to formal concepts, and a formal concept can be treated as an unordered sequence, we consider the Transformer to be proper for processing and capturing the information from bi-cliques. We have conducted experiments on several real-world datasets to show that after adjusting the architecture of the Transformer encoder, it can indeed capture pertinent information during training and leverage it for good link prediction performance.
Our main contributions are as summarized as follows:
@@
  • We introduce BicliqueEncoder, a novel method for bipartite link prediction using the Transformer encoder architecture. It can be applied to large datasets that the previous bi-cliques-based methods cannot handle, and its requirement of computational resources can be flexibly adjustable. We conduct experiments on five large datasets that exceed the capacity of traditional bi-clique-based approaches and five small datasets to demonstrate the efficacy and efficiency of our method.
@@ 2.1 Bipartite Networks and Bi-cliques
Bipartite Networks: A bipartite network $C$ is usually formalized as a bipartite graph a triple $(U,V,E)$, where $U$ and $V$ are two disjoint sets of nodes and $E\subseteq U\times V$ is a set of edges. An edge connecting a node $u\in U$ to another node $v\in V$ is denoted by $(u,v)$. Figure 1(a) illustrates an example of a bipartite graph $(U,V,E)$. In the graph all elements in $U$ are laid on the left-hand side and colored lime, and all elements in $V$ are laid on the right-hand side and colored blue. Note that every edges connects a node in $U$ and a node in $V$.
Subgraphs, Bi-cliques and Maximal Bi-cliques: A subgraph of a bipartite graph $G=(U,V,E)$ is a bipartite graph $G_{1}(U_{1},V_{1},E)$ such that $U_{1}\subseteq U$, $V_{1}\subseteq V$, $E_{1}\subseteq E\cap(U_{1}\times V_{1})$. The subgraph $G_{1}$ of is a bi-clique of $G$ if $E_{1}=U_{1}\times V_{1}$. We say that the bi-qlique $G_{1}$ is maximal if $G_{1}$ is not a subgraph of any other a maximal bi-clique of $G$, that is, for all $G_{2}=(U_{2},V_{2},E_{2})$ such that $U_{1}\subseteq U_{2}\subseteq U$ and $V_{1}\subseteq V_{2}\subseteq V$, $E_{2}=U_{2}\times V_{2}$ holds if and only if $U_{1}=U_{2}$ and $V_{1}=V_{2}$. Figure 1(b) gives an example of a maximal bi-clique.
@@ 2.2 Formal Concept Analysis (FCA)
Formal concept analysis (FCA) is a method for learning rules from a binary relational knowledge base, which is formal context. Given a formal context, the aim of FCA is to extract formal concepts and their hierarchical structure, which constitutes concept lattices. Here, we introduce the minimal necessary notions of FCA. For more detailed definitions and notions, readers are expected to refer to [20, 21, FCAsurvey2].
@@
A formal context is illustrated in a binary table, as exemplified in the left of Figure 2(a), where each row represents an object and each column represents an attribute. A cell in the table is marked with a cross if the object in its row possesses the attribute in its column.
Formal Concepts: In a context $\mathbb{K}=(G,M,I)$, for an object subset $A\subseteq G$ and an attribute subset $B\subseteq M$, $(A,B)$ is called a formal concept if for each $(A_{1},B_{1})$ such that $A\subseteq A_{1}\subseteq G$ and $B\subseteq B_{1}\subseteq M$, $A_{1}\times B_{1}\subseteq I$ is satisfied if and only if $A=A_{1}$ and $B=B_{1}$. If $(A,B)$ is a formal concept, $A$ is called an extent, and $B$ is called an intent.
Concept Lattices: Given a context $\mathbb{K}=(G,M,I)$, the concept lattice of context $\mathbb{K}$, denoted by $\underline{\mathfrak{B}}(\mathbb{K})$, is the structure that organizes the set of all concepts extracted from context $\mathbb{K}$ with the hierarchical order $<$. For two concepts $(A_{1},B_{1})$ and $(A_{2},B_{2})$, we write $(A_{1},B_{1})<(A_{2},B_{2})$ if $A_{1}\subset A_{2}$ (which mutually implies $B_{2}\subset B_{1}$[22]. A concept lattice is visualized with a line diagram. For example, the line diagram depicted in 2(b) represents the concept lattice of the context represented in 2(a). In the diagram, nodes represent formal concepts and lines represent hierarchical orders.
@@
Iceberg Concept Lattice: Given a context $\mathbb{K}=(G,M,I)$ and its concept lattice $\underline{\mathfrak{B}}(\mathbb{K})$, an iceberg concept lattice $\underline{\mathfrak{B}}^{\prime}(\mathbb{K})$ is defined to be the collection of all frequent concepts from the full concept lattice $\underline{\mathfrak{B}}(\mathbb{K})$. Here we say a concept is frequent if its extent has a size larger than a pre-determined threshold $\theta$ [^1] Since objects and attributes are dual, in some research, a frequent concept may also be dually defined for the concepts whose intent is larger than threshold $\theta$.. Formally, an iceberg concept lattice is defined to be $\{(A_{1},B_{1})\in\underline{\mathfrak{B}}(\mathbb{K})|\|A_{1}\|>\theta\}$. Figure 3 gives an example of iceberg concept lattices.
As introduced above, in a concept lattice, the concepts are ordered based on the sizes of their extents. Hence, we can easily derive that an iceberg concept lattice is formed of concepts from the top-most part of the full concept lattice, which is the reason it is named an iceberg concept lattice [18]. Since an iceberg concept lattices contains the concepts with the longest extents, it is often considered to represent the most “significant” part of the concept lattice, and is thus used in cases where the full concept lattice is considered redundant and needs to be reduced [18, 23].
Figure 3: Top-left: The upward iceberg concept lattice of the above example is extracted by considering extents with a length of at least 2, or intents with a length of at most 4. Top-right: The downward iceberg concept lattice of the above example is extracted by considering extents with a length of at most 5, or intents with a length of at least 3. Bottom: The intersection of two iceberg concept lattices.
@@ 2.4 Transformer Encoder
Transformer [19] is a deep learning network structure known for its ability to capture dependencies between elements within a sequence. A full Transformer architecture consists of an encoder and a decoder. The encoder processes the input sequence and outputs an “encoded” sequence, while the decoder recursively generates the output sequence based on the input sequence, the “encoded” sequence generated by the encoder, as well as the elements previously output by the decoder itself.
Usually, the decoder should be used together with the encoder in order to form the traditional sequence-to-sequence structure, which takes a sequence as input and outputs another sequence. However, both the decoder and the encoder can be used separately for different applications. For example, the Generative Pre-training Transformer (GPT) [24] is trained on a network with decoders only [24]. On the other hand, the Bidirectional Encoder Representations from Transformers (BERT) [25] model is trained on a network with encoders only [25]. Generally, since the decoder works recursively, it is preferred in scenes where the local order of the elements in the input/output sequences is considered important. The encoder, on the other hand, is preferred in scenes where the global dependencies of all elements in the input/output sequences are considered important. In this research, we plan to use the encoder of the Transformer to learn the information of bi-cliques.
@@
Figure 5: The detailed architecture of the standard Transformer encoder. Please refer to [19] for the original figure.
@@ 2.5 The Transformer Encoder and FCA
Given that the extents and intents of formal concepts are unordered sequences, there is no need to keep track of the order, and in our method, the positional encoding mechanism is removed. Details on our modifications to the Transformer encoder will be introduced in Section 4.
@@ 3 Related Work
Based on the structural hole theory [15] which has been mentioned above, several rule-based methods, such as Missbin and MF-NSS [12, 14, 13], have been proposed. These methods establish rules whereby if the overlap of two maximal bi-cliques exceeds a predefined threshold, the two bi-cliques merge into one, and the newly generated links are considered candidate links. These methods demonstrate better performance than some heuristic similarity-based methods such as Common Neighbors [7] and Jaccard [26].
However, manually crafted rules may not fully capture the information contained by bi-cliques. To better exploit the information of bi-cliques, researchers in [27] proposed FCA2VEC, a method to embed the nodes into a vector space based on their co-occurrence relationship within the maximal bi-cliques. The co-occurrences of nodes in a maximal bi-clique can be converted to the co-occurrences of objects in a formal concept, resembling the co-occurrences of words in a sentence. Thus, one can use the embedding models similar to the well-known word embedding model Word2Vec [28] to embed the nodes into vectors [27]. The FCA2VEC method presents a novel strategy for node embedding, by converting networks into formal contexts, extracting information from maximal bi-cliques using FCA, and subsequently processing the extracted information using embedding methods. Besides such a Word2Vec-like embedding method, in [16], the researchers proposed another method for node embedding based on bi-clique information using Bidirectional Long Short-Term Memory (Bi-LSTM) [29]. Both methods have shown good performances and proven that the information extracted from bi-cliques is useful for link prediction [16, 27]. However, these embedding methods cannot fully leverage all the information extractable by FCA, potentially limiting the prediction performance.
To address this limitation, researchers in [17] introduced BERT4FCA, a method utilizing BERT [25] to comprehensively learn information extracted by FCA, encompassing all formal concepts and their order relations, and leverage it for link prediction. BERT4FCA outperforms previous bi-clique-based approaches[17], underscoring the efficacy of incorporating richer information extracted by FCA for enhanced prediction performance. However, despite its remarkable prediction performance, it requires substantial memory and execution time. This problem renders it unsuitable for scenarios with limited computational resources.

### v2

@@ 1 Introduction
Link prediction in bipartite networks, or bipartite link prediction, is an important task for finding potential relations between two groups of entities. A bipartite network comprises two disjoint sets of entities and a set of edges, each connecting only two nodes from different sets. Two entities connected with an edge are considered related ([30, 1, 29]). Real-world bipartite networks often have missing or unobserved links, giving rise to the link prediction task for predicting the absence or presence of these links ([9, 19, 37]). Bipartite link prediction task has practical applications such as product recommendation in e-commerce (e.g., recommending products to customers based on co-purchases) ([15, 34]), chemical-disease interactions prediction in medical science to identify potential treatments ([21, 8, 39, 24]) and friend suggestions in social networking services (e.g., suggesting mutual friends on platforms like Facebook) ([1, 30, 37]).
@@
Most bipartite link prediction methods rely on structural features of the networks ([23]). Among all structural features, recent research highlights the importance of bi-cliques on bipartite link prediction ([23, 40, 38, 25]). A bi-clique is a complete sub-graph of a bipartite network where each entity in one set is connected to each entity in the other set. Bi-cliques can be regarded as clusters of strongly interconnected entities, representing cohesive groups within networks. For example, in a chemical-disease network, a bi-clique represents a group of similar chemicals and their affected diseases. Previous research has proposed the structural hole theory ([18]), suggesting that two strongly connected clusters sharing numerous nodes likely belong to a larger cluster, implying that current unlinked pairs within these clusters are likely to be connected. Thus, link prediction can be performed by extracting bi-cliques and analyzing their relationships. Various methods utilizing bi-cliques have been proposed and shown good performance in small datasets.
@@
However, all them face a severe problem: they are inapplicable to large-scale datasets, as they need to extract all bi-cliques from the network, which requires significant computational resources. Such an exhaustive enumeration is usually conducted by converting the network into an equivalent formal context and performing a full process of formal concept analysis (FCA) to extract all formal concepts, each corresponding to a maximal bi-clique in the original bipartite network, from the converted formal context. The time complexity of the full process of FCA is $O(C)$ where $C$ is the number of formal concepts, which is exponential to the number of entities and links in the network. Therefore, this bi-cliques extraction step may take a long time. Furthermore, after extraction, some methods ([22, 26]) require training complex neural-network models from all extracted bi-cliques, consuming extensive RAM and GPU memory. Consequently, these methods struggle with large and dense bipartite networks due to their high computational cost.
@@
To overcome this problem, we propose a novel and efficient method called BicliqueEncoder, which can be applied to large datasets with manageable and adjustable memory usage and execution time. The core feature of this method is that it does not extract all bi-cliques, but only extracts those that are considered “significant” in terms of their size. As introduced above, the process of extracting bi-cliques is equivalent to that of extracting formal concepts from a formal context converted from the original bipartite network. In the field of FCA, researchers have proposed the idea of iceberg concept lattice, which helps identify the possibility of only extracting the formal concepts of large sizes ([31]) without needing to exhaustively enumerate all formal concepts. Since large formal concepts correspond to large bi-cliques which contain more information on the relations between entities, we consider they might be more significant and contribute more to link prediction than smaller bi-cliques. Hence, if we only extract and process the bi-cliques corresponding to the formal concepts in the iceberg concept lattice, we may be able to capture the key information provided by the bi-cliques with a lot fewer computational resources.
@@
We adopt the Transformer encoder ([33]) to learn useful information and make link predictions because we believe that it can counterbalance the information loss caused by using only the extracted subset of all bi-cliques. Certainly, by applying the strategy that only extracts “significant” bi-cliques, we will leave out the information of all other “insignificant” bi-cliques in the bipartite network, which may cause a negative influence on the final prediction performance. For example, some nodes may not be included in the extracted bi-cliques. In such a case, the model cannot make predictions about such nodes when trained only on the extracted bi-cliques. To counterbalance such negative influence, we use the Transformer encoder to effectively process and capture the information from the original bipartite network as well as the extracted significant bi-cliques, and use the captured information to make link predictions. Transformer is a sequence-to-sequence deep learning framework adept at learning the dependencies within sequential data. Since bi-cliques are equivalent to formal concepts, and a formal concept can be treated as an unordered sequence (i.e. a set) we consider the Transformer to be proper for processing and capturing the information from bi-cliques. We have conducted experiments on several real-world datasets to show that after adjusting the architecture of the Transformer encoder, it can indeed capture pertinent information during training and leverage it for good link prediction performance.
Our main contributions are summarized as follows:
@@
  • We introduce BicliqueEncoder, a novel method for bipartite link prediction using the Transformer encoder architecture. It can be applied to large datasets that the previous bi-cliques-based methods cannot handle, and its requirement of computational resources is flexibly adjustable. We conduct experiments on five large datasets that exceed the capacity of traditional bi-clique-based approaches and five small datasets to demonstrate the efficacy and efficiency of our method.
@@ 2.1 Bipartite Networks and Bi-cliques
Bipartite Networks: A bipartite network $C$ is usually formalized as a bipartite graph, a triple $(U,V,E)$, where $U$ and $V$ are two disjoint sets of nodes and $E\subseteq U\times V$ is a set of edges. An edge connecting a node $u\in U$ to another node $v\in V$ is denoted by $(u,v)$. Figure 1(a) illustrates an example of a bipartite graph $(U,V,E)$. In the graph all elements in $U$ are laid on the left-hand side and colored lime, and all elements in $V$ are laid on the right-hand side and colored blue. Note that every edges connects a node in $U$ and a node in $V$.
Subgraphs, Bi-cliques and Maximal Bi-cliques: A subgraph of a bipartite graph $G=(U,V,E)$ is a bipartite graph $G_{1}(U_{1},V_{1},E)$ such that $U_{1}\subseteq U$, $V_{1}\subseteq V$, $E_{1}\subseteq E\cap(U_{1}\times V_{1})$. The subgraph $G_{1}$ of is a bi-clique of $G$ if $E_{1}=U_{1}\times V_{1}$. We say that the bi-clique $G_{1}$ is maximal if $G_{1}$ is not a subgraph of any other maximal bi-clique of $G$, that is, for all $G_{2}=(U_{2},V_{2},E_{2})$ such that $U_{1}\subseteq U_{2}\subseteq U$ and $V_{1}\subseteq V_{2}\subseteq V$, $E_{2}=U_{2}\times V_{2}$ holds if and only if $U_{1}=U_{2}$ and $V_{1}=V_{2}$. Figure 1(b) gives an example of a maximal bi-clique.
@@ 2.2 Formal Concept Analysis (FCA)
Formal concept analysis (FCA) is a method for learning rules from a binary relational knowledge base, which is formal context. Given a formal context, the aim of FCA is to extract formal concepts and their hierarchical structure, which constitutes concept lattices ([2, 13]).
@@
A formal context is illustrated in a binary table, as exemplified on the left of Figure 2(a), where each row represents an object and each column represents an attribute. A cell in the table is marked with a cross if the object in its row possesses the attribute in its column.
Formal Concepts: In a context $\mathbb{K}=(G,M,I)$, for an object subset $A\subseteq G$ and an attribute subset $B\subseteq M$, $(A,B)$ is called a formal concept if for each $(A_{1},B_{1})$ such that $A\subseteq A_{1}\subseteq G$ and $B\subseteq B_{1}\subseteq M$, $A_{1}\times B_{1}\subseteq I$ is satisfied if and only if $A=A_{1}$ and $B=B_{1}$. If $(A,B)$ is a formal concept, $A$ is called an extent, and $B$ is called an intent. In a formal concept, the extent contains all objects that share all attributes in the intent, and the intent contains all attributes that are shared in common by all objects in the extent.
Concept Lattices: Given a context $\mathbb{K}=(G,M,I)$, the concept lattice of context $\mathbb{K}$, denoted by $\underline{\mathfrak{B}}(\mathbb{K})$, is the structure that organizes the set of all concepts extracted from context $\mathbb{K}$ with the hierarchical order $<$. For two concepts $(A_{1},B_{1})$ and $(A_{2},B_{2})$, we write $(A_{1},B_{1})<(A_{2},B_{2})$ if $A_{1}\subset A_{2}$ (which mutually implies $B_{2}\subset B_{1}$) ([35]). A concept lattice is visualized with a line diagram. For example, the line diagram depicted in 2(b) represents the concept lattice of the context represented in 2(a). In the diagram, nodes represent formal concepts and lines represent hierarchical orders.
@@
Iceberg Concept Lattice: Given a context $\mathbb{K}=(G,M,I)$ and its concept lattice $\underline{\mathfrak{B}}(\mathbb{K})$, an iceberg concept lattice $\underline{\mathfrak{B}}^{\prime}(\mathbb{K})$ is defined to be the collection of all frequent concepts, which are frequent closed sets, from the full concept lattice $\underline{\mathfrak{B}}(\mathbb{K})$. Here we say a concept is frequent if its extent has a size larger than a pre-determined threshold $\theta$ [^1] Since objects and attributes are dual, in some research, a frequent concept may also be dually defined for the concepts whose intent is less than threshold $\theta$.. Formally, an iceberg concept lattice is defined to be $\{(A_{1},B_{1})\in\underline{\mathfrak{B}}(\mathbb{K})|\|A_{1}\|>\theta\}$. Figure 3 gives an example of iceberg concept lattices.
As introduced above, in a concept lattice, the concepts are ordered based on the sizes of their extents. Hence, we can easily derive that an iceberg concept lattice is formed of concepts from the top-most part of the full concept lattice, which is the reason it is named an iceberg concept lattice ([31]). Since an iceberg concept lattice contains the concepts with the largest extents, it is often considered to represent the most “significant” part of the concept lattice, and is thus used in cases where the full concept lattice is considered redundant and needs to be reduced ([31, 28]).
Figure 3: Top-left: The upward iceberg concept lattice of the above example is extracted by considering extents with a size of at least 2, or intents with a size of at most 4. Top-right: The downward iceberg concept lattice of the above example is extracted by considering extents with a size of at most 5, or intents with a size of at least 3. Bottom: The intersection of two iceberg concept lattices.
@@ 2.4 Transformer Encoder
Transformer ([33]) is a deep learning network structure known for its ability to capture dependencies between elements within a sequence. A full Transformer architecture consists of an encoder and a decoder. The encoder processes the input sequence and outputs an “encoded” sequence, while the decoder recursively generates the output sequence based on the input sequence, the “encoded” sequence generated by the encoder, as well as the elements previously output by the decoder itself.
Usually, the decoder should be used together with the encoder in order to form the traditional sequence-to-sequence structure, which takes a sequence as input and outputs another sequence. However, both the decoder and the encoder can be used separately for different applications. For example, the Generative Pre-training Transformer (GPT) ([27]) is trained on a network with decoders only ([27]). On the other hand, the Bidirectional Encoder Representations from Transformers (BERT) ([5]) model is trained on a network with encoders only ([5]). Generally, since the decoder works recursively, it is preferred in scenes where the local order of the elements in the input/output sequences is considered important. The encoder, on the other hand, is preferred in scenes where the global dependencies of all elements in the input/output sequences are considered important. In this research, we plan to use the encoder of the Transformer to learn the information of bi-cliques.
@@
Figure 5: The detailed architecture of the standard Transformer encoder. Please refer to ([33]) for the original figure.
@@ 2.5 The Transformer Encoder and FCA
Given that the extents and intents of formal concepts are unordered, there is no need to keep track of the order, and in our method, the positional encoding mechanism is removed. Details on our modifications to the Transformer encoder will be introduced in Section 4.
@@ 3 Related Work
Based on the structural hole theory ([18]) which has been mentioned above, several rule-based methods, such as Missbin and MF-NSS ([40, 25, 38]), have been proposed. These methods establish rules whereby if the overlap of two maximal bi-cliques exceeds a predefined threshold, the two bi-cliques merge into one, and the newly generated links are considered candidate links. These methods demonstrate better performance than some heuristic similarity-based methods such as Common Neighbors ([34]) and Jaccard ([16]).
However, manually crafted rules may not fully capture the information contained by bi-cliques. To better exploit the information of bi-cliques, researchers in ([7]) proposed FCA2VEC, a method to embed the nodes into a vector space based on their co-occurrence relationship within the maximal bi-cliques. The co-occurrences of nodes in a maximal bi-clique can be converted to the co-occurrences of objects in a formal concept, resembling the co-occurrences of words in a sentence. Thus, one can use the embedding models similar to the well-known word embedding model Word2Vec ([20]) to embed the nodes into vectors ([7]). The FCA2VEC method presents a novel strategy for node embedding, by converting networks into formal contexts, extracting information from maximal bi-cliques using FCA, and subsequently processing the extracted information using embedding methods. Besides such a Word2Vec-like embedding method, in ([22]), the researchers proposed another method for node embedding based on bi-clique information using Bidirectional Long Short-Term Memory (Bi-LSTM) ([10]). Both methods have shown good performances and proven that the information extracted from bi-cliques is useful for link prediction ([22, 7]). However, these embedding methods cannot fully leverage all the information extractable by FCA, potentially limiting the prediction performance.
To address this limitation, researchers in ([26]) introduced BERT4FCA, a method utilizing BERT ([5]) to comprehensively learn information extracted by FCA, encompassing all formal concepts and their order relations, and leverage it for link prediction. BERT4FCA outperforms previous bi-clique-based approaches [26]), underscoring the efficacy of incorporating richer information extracted by FCA for enhanced prediction performance. However, despite its remarkable prediction performance, it requires substantial memory and execution time. This problem renders it unsuitable for scenarios with limited computational resources.